OpenStudy (tumblewolf):

Trig Limit Help! 1) lim sin 2 z/z cos 3 z Z-0 2) lim tan 3x/tan 5x X-0 3) lim sin^2 theta/theta Theta-0 4) lim tan 2x/3x X-0

OpenStudy (tumblewolf):

@tkhunny do you have quick short cuts for these? I think 1= 2 but not sure

OpenStudy (tkhunny):

No shortcuts. There are a few things you can do. 1) Find a version of sin(W)/W which, when approaching zero is 1. Also, W/sin(W). 2) Is the same true of tan(W)/W? It would be helpful if it were so.

OpenStudy (anonymous):

\[\lim_{z \rightarrow 0}\frac{ \sin 2z }{ z \cos 3z }=2\lim_{z \rightarrow 0}\frac{ \sin 2z }{ 2z }*\lim_{z \rightarrow 0}\frac{ 1 }{ \cos 3z }=?\]

OpenStudy (anonymous):

\[\lim_{\theta \rightarrow 0}\frac{ \sin ^2\theta }{ \theta }=\lim_{\theta \rightarrow 0}\frac{ \sin ^2\theta }{ \theta^2 } \times \theta \] \[=\lim_{\theta \rightarrow 0}\left( \frac{ \sin \theta }{ \theta } \right)^2*\lim_{\theta \rightarrow 0}\theta =?\]

OpenStudy (anonymous):

i hope you can solve other .